Evaluate F(x)=∫ √t

2 + 9

x

4

.

a) F(4) b) F’(4) c) F’’(4)


1
Expert's answer
2021-12-14T02:51:47-0500

a)


"F(4)=\displaystyle\int_{4}^{4}\sqrt{t^2+9}dt=0"

b)

By the Fundamental Theorem of Calculus (part 1)


"F'(x)=\sqrt{x^2+9}"

"F'(4)=\sqrt{4^2+9}=5"

c)


"F''(x)=\dfrac{2x}{2\sqrt{x^2+9}}=\dfrac{x}{\sqrt{x^2+9}}"

"F''(4)=\dfrac{4}{\sqrt{4^2+9}}=\dfrac{4}{5}"